[figure: 9.7 cm]

LOW-TEMPERATURE TRANSPORT AND NOISE ANISOTROPY IN SmB_{6}

V.Glushkov^{a}, S.Demishev^{a}, M.Ignatov^{a}, N.Sluchanko^{a}, A.Savchenko^{b},

S.Safonov^{b}, V.Fillipov^{c}, Yu.Paderno^{c} aGeneral Physics Institute of RAS, Vavilov str., 38, Moscow, 119991, Russia

bPhysics Department, University of Exeter, Stocker Road, Exeter EX4 4QL, UK cInstitute for Problems of Materials Science, UA-252680, Kiev, Ukraine

Low temperature transport (Hall and Seebeck coefficients and resistivity) and noise characteristics have been measured in the model intermediate valence compound SmB_{6} along different crystallographic directions. It is shown that low-frequency resistance fluctuations are strongly anisotropic and the normalized electrical resistance noise demonstrates a low-temperature maximum similar to that one observed in the Hall mobility temperature dependence (N.E.Sluchanko et al., Phys.Rev.B 64 (2001), 153103; N.Sluchanko et al., Physica B (2002), in print). The anisotropy induced below 15K in this cubic narrow-gap semiconductor is discussed in terms of many-body states (exciton-polaron complexes) formation resulting from fast charge fluctuations on Sm- sites.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : V.V.Glushkov

Address for correspondence : General Physics Institute of RAS, 38, Vavilov str., Moscow, 119991, Russia

Email address : glushkov@lt.gpi.ru